# A Hitchhiker’s Guide to Privacy-Preserving Digital Payment Systems: A Survey on Anonymity, Confidentiality, and Auditability

### v2

@@ 3.5. Secure Multi-party Computation
Secure Multi-Party Computation (MPC) allows multiple parties to jointly compute a result over their private inputs without any party having to reveal its own input to the others. The key idea is that the parties collaborate to perform some computation and, at the end, they all learn the correct output of the function, but nothing else. This technique is used, e.g., in CoinParty (Ziegeldorf et al., 2018).
@@ 3.6. One-time Public Keys
Stealth addresses:
hide the recipient’s identity making each payment unlinkable. They are one-time public keys generated by the transaction sender with public key provided by the receiver. Recipient does not update his public key, but the sender can derive fresh one-time keys for each transaction. Let $\mathrm{G}$ be a generator of an elliptic curve group $\mathbb{G}$ with scalar field $\mathbb{Z}_{q}$. Each receiver $R$ has a public view key $V=vG$ and a public spend key $S=sG$, where $v,s\in\mathbb{Z}_{q}$ are secret scalars. A sender chooses a random scalar $r\in\mathbb{Z}_{q}$ and computes the shared secret $H=\mathcal{H}_{p}(rV)\in\mathbb{Z}_{q}$, with $\mathcal{H}_{p}(\cdot)$ a hash-to-scalar function, and the one-time public key $P=HG+S$. The sender includes $P$ and $R=rG$ in the transaction. The recipient can scan the ledger and for each $P$ and $R$ computes $P^{\prime}=\mathcal{H}_{p}(vR)G+S$. If $P^{\prime}=P$ then the output belongs to the recipient.
Updatable public keys:
allow to update the recipient’s public key in a decentralized manner. Quisquis (Fauzi et al., 2019) introduces this technique to update both sender and recipient public keys. Given a prime-order group $\mathbb{G}_{p}$ with $p$ the prime order of the group, and $G$ its generator, each user has a secret key $s_{i}$ and a public key $P_{i}=(G^{r},G^{r\cdot s_{i}})=(U,V)$, for some random value $r\in\mathbb{Z}_{q}$. To perform a transaction, the sender updates the public key by choosing a random value $z\in\mathbb{Z}_{q}$ and computing $P^{\prime}_{i}=(U^{z},V^{z})$. The updated key is indistinguishable from a fresh key. The recipient verifies whether a given updated public key $P^{\prime}_{i}$ corresponds to his secret key $s_{i}$. This works because updated keys still verify correctly with the original secret key.
Figure 1. Dimension of analysis along which we review privacy-preserving digital payment systems.:
@@ 4. Design Dimensions of Privacy-Preserving Digital Payment Systems
The research community has explored various solutions to ensure anonymity, confidentiality, and auditability in crypto-assets and digital payment systems. We identify key dimensions of analysis—protocol, value representation, privacy features, and auditability—and highlight the most significant design choices. These dimensions are interconnected and jointly embed privacy features into the system. Figure 1 outlines the dimensions that structure this section. Table 1 describes the labels used in Table LABEL:tab:sota-cryptocurrencies, which summarizes our literature review.
@@ 4.1. Protocol
Table 1. Description of the labels used in Table LABEL:tab:sota-cryptocurrencies.
Dimension Legend Protocol Type CD CBDC D Digital payment system Blockchain Compatibility B Bitcoin E Ethereum Value Representation Data Model A Account-based CT Chaumian token T2B Transaction per bank table UT Unspent token Other Auditability AB Anonymity budget L Operating limits RA Revocable anonymity TR Coin tracing ZD Verification without disclosure Cryptographic Primitives BS Blind signature C Commitment DS Digital signature GS Group signatures HC Homomorphic commitment IBE Identity-based encryption MS Multi signature MT Merkle tree O-RAM Oblivious RAM PE Public key encryption RS Ring signature SA Stealth address UP Updatable public keys ZK Zero-knowledge proof $X^{t}$ Threshold-based version of $X$ $X^{r}$ Randomizable version of $X$ XX* Custom variant of $X$ Generics Yes $\times$ No Partial $\Leftrightarrow$ Alternative Abstracted
@@ 4.1.1. Generations of Privacy-preserving Digital Payment Systems
Privacy-preserving digital payment systems have evolved through three generations, shaped by blockchain adoption, advances in cryptography, and interest in CBDCs. The first generation focuses on centralized authorities, emphasizing sender anonymity. The second, emerging with Bitcoin, enables decentralized value transfer and focuses on unlinkability and untraceability. The third adds mechanisms to balance privacy with auditability and regulatory compliance.
@@
The first generation focuses on the design of privacy-preserving value transfer systems relying on digital signatures and a centralized authority (i.e., the issuing bank) (Chaum, 1983). The fundamental problem relates to guaranteeing anonymity of sender (i.e., the payer) although he must interact with the central bank to obtain valid tokens (representing coins) to perform subsequent payments. Chaum (Chaum, 1983) introduces one of the first untraceable payment systems, commonly known as ecash. Later, several works improved ecash with different features (Andola et al., 2021), including support for different coin denominations (e.g., (Dold, 2019)), partial spending (e.g., (Dold, 2019)), token transferability (e.g., (Hayes, 1990; Sai Anand and Madhavan, 2000; Baldimtsi et al., 2015; Bauer et al., 2021; Chase et al., 2014)), offline spending (e.g., (Chaum et al., 1990)), and auditability (e.g., (Camenisch et al., 2006; Dold, 2019)).
@@
The second generation emerged with the introduction of Bitcoin (Nakamoto, 2008). Bitcoin and, later, Ethereum (Buterin et al., 2013) propose public and permissionless blockchains, meaning that the ledger contains transactions visible to everyone and anyone can participate in the process of updating the public ledger. Being public, these blockchains use addresses to exchange assets in a pseudo-anonymous manner, as addresses are not directly tied to user identities. Public blockchains cannot easily and fully satisfy two privacy-related features: untraceability and unlinkability (Rahalkar and Virgaonkar, 2024). Untraceability ensures that, for any given incoming transaction, all potential senders are equally likely. Unlinkability means that, for any two outgoing transactions, there is no way to prove they were sent to the same recipient. While generating a new address for each transaction can enhance these properties, various studies indicate that achieving true anonymity and unlinkability on public blockchains remains challenging. An analysis of the transaction graph can reveal clusters of pseudonyms or patterns that link with high probability two or more addresses (e.g., (Meiklejohn et al., 2016; Möser et al., 2018; Kappos et al., 2018; Herskind et al., 2020)). The second generation of solutions mainly focuses on improving unlinkability, anonymity, and confidentiality, often considering decentralized architectures (e.g., (Miers et al., 2013; Van Saberhagen, 2013; Fauzi et al., 2019; Baudet et al., 2023; Friolo et al., 2024; Sarencheh et al., 2025; Androulaki et al., 2013; Herrera-Joancomartí, 2015; Meiklejohn et al., 2013)).
@@
We classify this second generation of approaches into three main categories: privacy-native protocols, mixing protocols, and off-ledger solutions. Changing the protocol of existing blockchains is hard, as it involves forks. Therefore, privacy-native protocols (mostly) defines novel solutions with improved confidentiality, anonymity, and unlinkability features (e.g., (Van Saberhagen, 2013; Noether et al., 2016; Ben Sasson et al., 2014; Tomescu et al., 2022)). Conversely, mixing protocols propose supplementary services for permissionless blockchains aimed to collect a group of transactions, scramble senders and receivers, and publish them as a single group transaction (e.g., (Maxwell, 2013; Bissias et al., 2014; Bonneau et al., 2014; Valenta and Rowan, 2015; Ruffing and Moreno-Sanchez, 2017; Ziegeldorf et al., 2018; Qin et al., 2023; Wang et al., 2023)). Mixing protocols break the link between sender and receiver, thus reducing the possibility of linking together addresses of the blockchain. Off-ledger solutions are a specific approach related to blockchains (that serve as Layer 1) and the ability to build decentralized peer-to-peer Layer 2 systems on top of them. Layer 2 solutions process transactions off the main chain, reducing the need to record every transaction on the shared ledger (e.g., (Qin et al., 2023; Wang et al., 2023; Baudet et al., 2023; Seres et al., 2019)). Since we focus on on-ledger privacy, we omit a detailed analysis of Layer 2 privacy solutions.
@@ 4.1.2. System Architecture
Architecture of Mixing Protocols
Both centralized (e.g., (Bonneau et al., 2014; Valenta and Rowan, 2015; Heilman et al., 2016; Qin et al., 2023; Liu et al., 2018)) and decentralized solutions (e.g., (Ruffing et al., 2014; Ziegeldorf et al., 2018; Bissias et al., 2014; Ruffing and Moreno-Sanchez, 2017; Nosouhi et al., 2023; Wang et al., 2023)) have been proposed to implement mixing protocols. In centralized solutions, a third party (mixer or tumbler) provides the mixing service for a fee. While such designs are simple and support large anonymity sets, they introduce risks such as availability, theft, and potential privacy breaches. Therefore, different trust models have been proposed to mitigate these concerns (see Section 4.1.4). Decentralized mixers remove the need for a trusted party by using cryptographic protocols or peer-to-peer coordination, enhancing security and censorship resistance, but they often incur higher complexity, require user coordination, and typically support smaller anonymity sets (e.g., Xim (Bissias et al., 2014), CoinShuffle++ (Ruffing et al., 2016), and ValueShuffle (Ruffing and Moreno-Sanchez, 2017)). Both centralized and decentralized approaches face challenges such as Sybil attacks [^4] In a Sybil attack, an adversary injects fake participants into the mix to degrade anonymity. and regulatory scrutiny.
@@ 4.1.3. Protocol Integration and Compatibility with Existing Blockchains
Mixing protocols propose a supplementary services for existing blockchains (mostly for Bitcoin), and are specifically designed to be fully compatible (e.g., (Ruffing et al., 2016; Valenta and Rowan, 2015; Qin et al., 2023; Ruffing and Moreno-Sanchez, 2017)). Another category of compatible solutions leverages blockchain programmability to introduce privacy-preserving overlays (e.g., (Bünz et al., 2020; Baudet et al., 2023)). For instance, Zether (Bünz et al., 2020) uses a smart contract on Ethereum to store and manage private accounts; Zef (Baudet et al., 2023) introduces opaque coins in the form of off-chain certificates exchanged between Ethereum user accounts.
@@
Most privacy-enhancing proposals are non-compatible with the blockchain protocol they extend. For example, Zerocoin (Miers et al., 2013) is an overlay over Bitcoin that allows user to convert bitcoins in zerocoins. A zerocoin can be proved (in ZK) to originate from a valid and unspent bitcoin, without explicitly revealing such link. Zerocoin does not provide full anonymity: it obscures coin traceability, but it does not hide the number of transactions, recipient addresses, and balances. Building on Zerocoin, Zerocash (Ben Sasson et al., 2014) enhances privacy by concealing both sender and recipient addresses, supporting arbitrary transaction values, and enabling change handling. Both Zerocoin and Zerocash require new transaction types and payment semantics, and break compatibility with Bitcoin. Garman et al. (Garman et al., 2017) extends Zerocash with no backward compatibility, by introducing policy enforcement and tracing of users and tainted coins—features that further diverge from Bitcoin’s existing design. Yu et al. (Yu et al., 2024) propose an overlay on Bitcoin for anonymous yet regulated transactions using a coin backed by bitcoins. Their approach relies on three centralized entities that cooperate to enable batch linkability and payer tracing while preserving privacy. These changes are not backward-compatible with Bitcoin.
@@
Most approaches present greenfield solutions for privacy-preserving payments or asset exchanges (e.g., (Van Saberhagen, 2013; Poelstra, 2016; Noether et al., 2016; Narula et al., 2018; Fauzi et al., 2019; Wüst et al., 2019; Zhang et al., 2019; Androulaki et al., 2020; Chen et al., 2020)), often extending open-source solutions (e.g., (Li et al., 2021; Papadoulis et al., 2024)). Both centralized (e.g., (Cecchetti et al., 2017; Zhang et al., 2019)) and decentralized architectures (e.g., (Narula et al., 2018; Androulaki et al., 2020; Lin et al., 2020; Ma et al., 2021; Goodell et al., 2023; Elfadul et al., 2024)) exist, with a sensible preference for the latter. Early approaches like CryptoNote (Van Saberhagen, 2013), MimbleWimble (Poelstra, 2016), and Monero (Noether et al., 2016) address the Bitcoin’s privacy limitations by still proposing fully decentralized alternatives. CryptoNote (Van Saberhagen, 2013) enhances sender and receiver anonymity through one-time ring signatures and NIZK proofs, inspiring solutions like ByteCoin, DigitalNote, and Aeon (Amarasinghe et al., 2019). MimbleWimble (Poelstra, 2016) improves privacy and scalability by aggregating confidential transactions. Grin [^5] https://grin.mw/ and Beam [^6] https://www.beam.mw/ build on MimbleWimble ideas. Monero (Noether et al., 2016) strengthens privacy with sender, receiver, and network anonymity as well as confidential transactions.
Other approaches propose partially decentralized solutions (e.g., (Wüst et al., 2019)), combine concepts from centralized mixers (e.g., (Fauzi et al., 2019)), or consider hierarchical bank-intermediated settings (e.g., (Cecchetti et al., 2017; Narula et al., 2018)). Wüst et al. (Wüst et al., 2019) propose PRCash to strike a trade-off between fast payments, user privacy, and regulatory oversight: it includes two centralized entities for issuance and oversight and a permissioned blockchain, with MimbleWimble-like transactions. Quisquis (Fauzi et al., 2019) builds on the concept of mixing to redistribute wealth: each transaction updates a bunch of other accounts on chain, thus preventing addresses from appearing multiple times on the ledger, reducing the risk of de-anonymization and linkability. Cecchetti et al. (Cecchetti et al., 2017) propose Solidus, a hierarchical, bank-intermediated systems where a small number of banks manage transactions of on-chain assets on behalf of a large number of users; this setup ensures transaction-graph confidentiality. [^7] While banks can identify their respective clients, other entities only learn the identities of the banks.

### v3

@@ 3.5. Secure Multi-party Computation
Secure Multi-Party Computation (MPC) allows multiple parties to jointly compute a result over their private inputs without any party having to reveal its own input to the others. The key idea is that the parties collaborate to perform some computation and, at the end, they all learn the correct output of the function, but nothing else. Beyond applications such as CoinParty (Ziegeldorf et al., 2018), MPC is also used to decentralize trusted setup procedures in zk-SNARKs, enabling security to hold as long as at least one participant behaves honestly.
@@ 3.6. One-time Public Keys
Stealth addresses
They hide the recipient’s identity making each payment unlinkable. They are one-time public keys generated by the transaction sender with public key provided by the receiver. Recipient does not update his public key, but the sender can derive fresh one-time keys for each transaction. Let $\mathrm{G}$ be a generator of an elliptic curve group $\mathbb{G}$ with scalar field $\mathbb{Z}_{q}$. Each receiver $R$ has a public view key $V=vG$ and a public spend key $S=sG$, where $v,s\in\mathbb{Z}_{q}$ are secret scalars. A sender chooses a random scalar $r\in\mathbb{Z}_{q}$ and computes the shared secret $H=\mathcal{H}_{p}(rV)\in\mathbb{Z}_{q}$, with $\mathcal{H}_{p}(\cdot)$ a hash-to-scalar function, and the one-time public key $P=HG+S$. The sender includes $P$ and $R=rG$ in the transaction. The recipient can scan the ledger and for each $P$ and $R$ computes $P^{\prime}=\mathcal{H}_{p}(vR)G+S$. If $P^{\prime}=P$ then the output belongs to the recipient.
Updatable public keys
They allow to update the recipient’s public key in a decentralized manner. Quisquis (Fauzi et al., 2019) introduces this technique to update both sender and recipient public keys. Given a prime-order group $\mathbb{G}_{p}$ with $p$ the prime order of the group, and $G$ its generator, each user has a secret key $s_{i}$ and a public key $P_{i}=(G^{r},G^{r\cdot s_{i}})=(U,V)$, for some random value $r\in\mathbb{Z}_{q}$. To perform a transaction, the sender updates the public key by choosing a random value $z\in\mathbb{Z}_{q}$ and computing $P^{\prime}_{i}=(U^{z},V^{z})$. The updated key is indistinguishable from a fresh key, thanks to the Decisional Diffie-Hellman assumption. The recipient verifies whether a given updated public key $P^{\prime}_{i}$ corresponds to his secret key $s_{i}$. This works because updated keys still verify correctly with the original secret key.
Figure 1. Dimension of analysis along which we review privacy-preserving digital payment systems.
Dimension of analysis along which we review privacy-preserving digital payment systems.
Table 1. Labels and notation used in Table LABEL:tab:sota-cryptocurrencies.
Dimension Legend Protocol Type CD CBDC D Digital payment system Blockchain Compatibility B Bitcoin E Ethereum Value Representation Data Model A Account-based CT Chaumian token T2B Transaction per bank table UT Unspent token Other Auditability AB Anonymity budget L Operating limits RA Revocable anonymity TR Coin tracing ZD Verification without disclosure Cryptographic Primitives BS Blind signature C Commitment DS Digital signature GS Group signatures HC Homomorphic commitment IBE Identity-based encryption MS Multi signature MT Merkle tree O-RAM Oblivious RAM PE Public key encryption RS Ring signature SA Stealth address UP Updatable public keys ZK Zero-knowledge proof $X^{t}$ Threshold-based version of $X$ $X^{r}$ Randomizable version of $X$ XX* Custom variant of $X$ Notation Yes $\times$ No Partial $\Leftrightarrow$ Alternative Abstracted
@@ 4. Design Dimensions of Privacy-Preserving Digital Payment Systems
The research community has explored various solutions to ensure anonymity, confidentiality, and auditability in crypto-assets and digital payment systems. We identify key dimensions of analysis—protocol, value representation, privacy features, and auditability—and highlight the most significant design choices. [^4] To avoid terminological ambiguity, we align with De Roode and Everts (de Roode and Everts, 2025), who distinguish privacy (a user right) from privacy-preservation (a system property) and define anonymity, unlinkability, and auditability in a unified DLT terminology. These dimensions are interconnected and jointly embed privacy features into the system. Figure 1 outlines the dimensions that structure this section. Table 1 describes the labels used in Table LABEL:tab:sota-cryptocurrencies, which summarizes our literature review. [^5] Some properties shown in Figure 1 and Table LABEL:tab:sota-cryptocurrencies (e.g., double-spending protection, ledger state updates) are treated as supporting mechanisms underpinning correctness, privacy, or auditability, and are discussed within those dimensions rather than as independent classification axes.
@@ 4.1.1. Generations of Privacy-preserving Digital Payment Systems
We introduce the notion of generations of privacy-preserving digital payment systems to group works that share similar assumptions, deployment contexts, and design goals. Under this lens, first-generation systems typically rely on centralized issuers and focus on payer anonymity; second-generation systems emphasize decentralized value transfer, unlinkability, and untraceability; while third-generation systems balance privacy with auditability and regulatory compliance. [^6] These groupings are not strictly chronological, and several systems naturally exhibit hybrid characteristics across generations. The “generations” serve as conceptual lens to interpret recurring design goals and assumptions across the dimensions discussed in this section.
@@
The first generation of privacy-preserving digital payment systems focuses on centralized value transfer architectures, where a trusted issuing authority (typically a bank) is responsible for minting and redeeming digital coins, while cryptographic mechanisms are used to protect the payer’s anonymity during spending. The central challenge in this setting is reconciling the need for user anonymity with the requirement that the issuer validate, redeem, and potentially audit issued tokens. Chaum (Chaum, 1983) introduces the seminal untraceable payment system, known as ecash [^7] Later, Chaum et al. (Chaum et al., 2021) show how these mechanisms can be embedded into a two-tier, regulator-compliant retail CBDC architecture that preserves cash-like payer privacy while enforcing income transparency and AML/CFT through commercial banks., which employs blind signatures to prevent the issuing bank from linking withdrawn coins to subsequent payments (see Section 4.2 for details). Building on this foundation, a large body of work explored extensions of ecash (Andola et al., 2021), including different coin denominations (e.g., (Dold, 2019)), partial spending (e.g., (Dold, 2019; Okamoto and Ohta, 1992)), token transferability (e.g., (Hayes, 1990; Sai Anand and Madhavan, 2000; Baldimtsi et al., 2015; Bauer et al., 2021; Chase et al., 2014)), offline spending (e.g., (Chaum et al., 1990)), and controlled forms of auditability (e.g., (Camenisch et al., 2006; Schoenmakers, 1998; Brands, 1994; Camenisch and Lysyanskaya, 2001; Batten and Yi, 2019)). [^8] In contrast to Chaum’s ecash, Netcoin (Tewari and Nuallain, 2015) proposes a fiat-backed mint-based system in which coins are bound to user identities and fully traceable (no anonymity). This highlights how privacy was not a universal requirement in early centralized ecash designs.
@@
The second generation emerged with the introduction of Bitcoin (Nakamoto, 2008). Bitcoin and, later, Ethereum (Buterin et al., 2013) propose public and permissionless blockchains, meaning that the ledger contains transactions visible to everyone and anyone can participate in the process of updating the public ledger. Being public, these blockchains use addresses to exchange assets in a pseudo-anonymous manner, as addresses are not directly tied to user identities. [^9] Bitcoin addresses are designed to be ephemeral (though often reused), whereas Ethereum relies on persistent pseudonymous accounts. Public blockchains cannot easily and fully satisfy two privacy-related features: untraceability and unlinkability (Rahalkar and Virgaonkar, 2024). Untraceability ensures that, for any given incoming transaction, all potential senders appear equally likely. Unlinkability means that, for any two outgoing transactions, there is no way to prove they were sent to the same recipient. Untraceability and unlinkability can arise at different layers: at the ledger level, they refer to preventing correlations between on-chain records; at the network level, they refer to preventing correlations from communication patterns [^10] Unless otherwise stated, we primarily focus on ledger-level privacy, and discuss network-layer aspects in Section 4.3.4.. While generating a new address for each transaction can enhance these properties, various studies indicate that achieving true anonymity and unlinkability on public blockchains remains challenging. An analysis of the transaction graph can reveal clusters of pseudonyms or patterns that link with high probability two or more addresses (e.g., (Meiklejohn et al., 2013; Möser et al., 2018; Kappos et al., 2018; Herskind et al., 2020)). The second generation of solutions mainly focuses on improving unlinkability, anonymity, and confidentiality, often considering decentralized architectures (e.g., (Miers et al., 2013; Van Saberhagen, 2013; Fauzi et al., 2019; Baudet et al., 2023; Friolo et al., 2026; Sarencheh et al., 2025; Androulaki et al., 2013; Herrera-Joancomartí, 2015; Meiklejohn et al., 2013)).
@@
We classify this second generation of approaches into three main categories: privacy-native protocols, mixing protocols, and off-ledger solutions. Changing the protocol of existing blockchains is hard, as it involves forks. Therefore, privacy-native protocols (mostly) defines novel solutions with improved confidentiality, anonymity, and unlinkability features (e.g., (Van Saberhagen, 2013; Noether et al., 2016; Ben Sasson et al., 2014; Tomescu et al., 2022)). Conversely, mixing protocols propose supplementary services for permissionless blockchains aimed to collect a group of transactions, scramble senders and receivers, and publish them as a single group transaction (e.g., (Maxwell, 2013; Bissias et al., 2014; Bonneau et al., 2014; Valenta and Rowan, 2015; Ruffing and Moreno-Sanchez, 2017; Ziegeldorf et al., 2018; Qin et al., 2023; Wang et al., 2023)). Mixing protocols break the link between sender and receiver, thus reducing the possibility of linking together addresses of the blockchain (an analysis of protocols in Appendix A.3.2). Off-ledger solutions are a specific approach related to blockchains (that serve as Layer 1) and the ability to build decentralized peer-to-peer Layer 2 systems on top of them. Layer 2 solutions process transactions off the main chain, reducing the need to record every transaction on the shared ledger (e.g., (Qin et al., 2023; Wang et al., 2023; Baudet et al., 2023; Seres et al., 2019)). Focusing on on-ledger privacy, we omit the analysis of Layer 2 privacy solutions.
@@ 4.1.3. Protocol Integration and Compatibility with Existing Blockchains
Mixing protocols propose supplementary services for existing blockchains (mostly for Bitcoin), and are specifically designed to be fully compatible (e.g., (Ruffing et al., 2016; Valenta and Rowan, 2015; Qin et al., 2023; Ruffing and Moreno-Sanchez, 2017)). Another category of compatible solutions leverages blockchain programmability to introduce privacy-preserving overlays (e.g., (Bünz et al., 2020; Baudet et al., 2023)). For instance, Zether (Bünz et al., 2020) uses a smart contract on Ethereum to store and manage private accounts; Zef (Baudet et al., 2023) introduces opaque coins in the form of off-chain certificates exchanged between Ethereum user accounts.
@@
Most privacy-enhancing proposals are non-compatible with the blockchain protocol they extend. For example, Zerocoin (Miers et al., 2013) is an overlay over Bitcoin that allows user to convert bitcoins in zerocoins. A zerocoin can be proved (in ZK) to originate from a valid and unspent bitcoin, without explicitly revealing such link. Zerocoin does not provide full anonymity: it obscures coin traceability, but it does not hide the number of transactions, recipient addresses, and balances. Building on Zerocoin, Zerocash (Ben Sasson et al., 2014) enhances privacy by concealing both sender and recipient addresses, supporting arbitrary transaction values, and enabling change handling. Both Zerocoin and Zerocash require new transaction types and payment semantics, and break compatibility with Bitcoin. While, in theory, they could be integrated if a super-majority of miners and nodes adopt the new protocol, the authors do not expect such integration and instead recommend deploying Zerocash as a separate system. Garman et al. (Garman et al., 2017) extends Zerocash with no backward compatibility, by introducing policy enforcement and tracing of users and tainted coins—features that further diverge from Bitcoin’s existing design. Yu et al. (Yu et al., 2024) propose an overlay on Bitcoin for anonymous yet regulated transactions using a coin backed by bitcoins. Their approach relies on three centralized entities (i.e., bank, tracing authority, and technical support) that cooperate to enable batch linkability and payer tracing while preserving privacy. These changes are not backward-compatible with Bitcoin.
@@
Most approaches present greenfield solutions for privacy-preserving payments or asset exchanges (e.g., (Van Saberhagen, 2013; Poelstra, 2016; Noether et al., 2016; Narula et al., 2018; Fauzi et al., 2019; Wüst et al., 2019; Zhang et al., 2019; Androulaki et al., 2020; Chen et al., 2020)), often extending open-source solutions (e.g., (Li et al., 2021; Papadoulis et al., 2024)). Both centralized (e.g., (Cecchetti et al., 2017; Zhang et al., 2019)) and decentralized architectures (e.g., (Narula et al., 2018; Androulaki et al., 2020; Lin et al., 2020; Ma et al., 2021; Goodell et al., 2023; Elfadul et al., 2024)) exist, with a sensible preference for the latter. Early approaches like CryptoNote (Van Saberhagen, 2013), MimbleWimble (Poelstra, 2016), and Monero (Noether et al., 2016) address Bitcoin’s privacy limitations by still proposing fully decentralized alternatives. CryptoNote (Van Saberhagen, 2013) enhances sender and receiver anonymity through one-time ring signatures and NIZK proofs, inspiring solutions like ByteCoin, DigitalNote, and Aeon (Amarasinghe et al., 2019). MimbleWimble (Poelstra, 2016) improves privacy and scalability by aggregating confidential transactions. However, this approach does not guarantee unlinkability among participants who undertake the aggregation. Grin [^12] https://grin.mw/ and Beam [^13] https://www.beam.mw/ build on MimbleWimble ideas. Monero (Noether et al., 2016) strengthens privacy with sender, receiver, and network anonymity as well as confidential transactions. Despite its strong anonymity measures, research has identified vulnerabilities in Monero’s mixing selection strategy that could lead to transaction de-anonymization (e.g., (Möser et al., 2018; Kappos et al., 2018)).
Other approaches propose partially decentralized solutions (e.g., (Wüst et al., 2019)), combine concepts from centralized mixers (e.g., (Fauzi et al., 2019)), or consider hierarchical bank-intermediated settings (e.g., (Cecchetti et al., 2017; Narula et al., 2018)). Wüst et al. (Wüst et al., 2019) propose PRCash to strike a trade-off between fast payments, user privacy, and regulatory oversight: it includes two centralized entities for issuance and oversight and a permissioned blockchain, with MimbleWimble-like transactions (i.e., commitment-based; see Section 4.2.2). Quisquis (Fauzi et al., 2019) builds on the concept of mixing to redistribute wealth: each transaction updates a bunch of other accounts on chain, thus preventing addresses from appearing multiple times on the ledger, reducing the risk of de-anonymization and linkability. Quisquis ensures integrity through ZKPs while maintaining privacy using commitments and updatable public keys.